# Between Lines of Code: Unraveling the Distinct Patterns of Machine and Human Programmers

### v3

@@ Abstract
Large language models have catalyzed an unprecedented wave in code generation. While achieving significant advances, they blur the distinctions between machine- and human-authored source code, causing integrity and authenticity issues of software artifacts. Previous methods such as DetectGPT have proven effective in discerning machine-generated texts, but they do not identify and harness the unique patterns of machine-generated code. Thus, its applicability falters when applied to code. In this paper, we carefully study the specific patterns that characterize machine- and human-authored code. Through a rigorous analysis of code attributes such as lexical diversity, conciseness, and naturalness, we expose unique patterns inherent to each source. We particularly notice that the syntactic segmentation of code is a critical factor in identifying its provenance. Based on our findings, we propose DetectCodeGPT, a novel method for detecting machine-generated code, which improves DetectGPT by capturing the distinct stylized patterns of code. Diverging from conventional techniques that depend on external LLMs for perturbations, DetectCodeGPT perturbs the code corpus by strategically inserting spaces and newlines, ensuring both efficacy and efficiency. Experiment results show that our approach significantly outperforms state-of-the-art techniques in detecting machine-generated code  [^1] Our code and data are available at https://github.com/YerbaPage/DetectCodeGPT.
@@ III-A1 Lexical Diversity
keyword
def, return, else, if, for, while, $\ldots$
identifier
identifier, type_identifier
literal
string_content, integer, true, false, $\ldots$
operator
<, >, =, ==, +, $\ldots$
syntactic symbol
comment
comment
whitespace
@@ III-B Experimental Setup
We choose the state-of-the-art CodeLlama model [6] to generate code. Limited by our computational resources, we use the version with 7B parameters. As for the decoding strategies, we adopt the top-$p$ sampling method [36] with $p$=$0.95$ following [1]. The temperature $T$ is an important parameter controlling the diversity of the generated code [36]. Since current LLMs on code are usually evaluated across different decoding temperatures [5, 37, 1, 6], we generate code with $T=0.2$ and $T=1.0$ to capture the model’s behavior under different settings. The maximum length of the generated code is set to 512 tokens. All experiments are conducted on 2 NVIDIA RTX 4090 GPUs with 24GB memory.
@@ III-D2 Syntax Element Distribution
The analysis of syntax element distributions, as visualized in Figure 1, reveals intriguing insights into the coding conventions and stylistic nuances between human- and machine-authored code. The “keyword”, “operator”, “syntactic symbols”, and “whitespace” proportions remain largely consistent between human and machine-authored code, across both temperatures. This consistency suggests that the foundational syntactical elements manifest similarly in both datasets. Delving deeper into the more nuanced discrepancies, a few categories emerge that underscore the differential preferences or tendencies of human and machine writers (statistical significance $p<0.01$):
@@ III-D3 Zipf’s and Heaps’ Laws
Figure 2 offers a comprehensive understanding of coding tendencies of both human and machine programmers. Starting with Zipf’s law, Figure 1(a) and 1(b) both delineate similar trends for human and machine programmers, corroborating the law’s applicability. And we can observe machine’s heightened proclivity towards tokens ranked between 10 and 100, especially at $T$=$0.2$. Turning our attention to Heaps’ law, the near-linear trends in Figure 1(c)-1(d) reaffirm the law’s validity. Also, there’s a noticeable shallowness in the slope for machine’s code at $T$=$0.2$ revealing machine’s decreased lexical diversity.
@@ III-D5 Token Likelihood and Rank
ALL -1.658 -0.827 0.831 0.319 0.811 0.492
Table III summarises the comparison results in terms of each token category at $T=0.2$. An intriguing finding is that whitespace tokens stand out with the highest deviation of naturalness, surpassing even the combined use of all tokens. This highlights a distinctive aspect of coding styles: machines, trained on extensive datasets, typically generate code with regular, predictable whitespace patterns. Humans, however, influenced by individual styles and practices, exhibit a wider variety in their use of whitespaces. The distinct patterns in machine-generated whitespaces, therefore, point to an inherent variation in coding style between machine and human.
@@ IV-C1 Space Insertion
Let $C$ represent the set of all possible locations to insert spaces in a code segment. We randomly select a subset $C_{s}\subseteq C$ such that $|C_{s}|=\alpha\times|C|$, where $\alpha\in[0,1]$ is a fraction representing the code locations. For each location $c\in C_{s}$, we introduce a variable number of spaces, $n_{\text{spaces}}(c)$, which is drawn from a Poisson distribution $\mathcal{P}(\lambda_{\text{spaces}})$. Mathematically, this can be represented as:
@@ IV-C2 Newline Insertion
We split the code into lines and obtain a set $L$ of lines. A subset $L_{n}\subseteq L$ is then chosen randomly, where $|L_{n}|=\beta\times|L|$, with $\beta\in[0,1]$ denoting the proportion of the line locations. For each line $l\in L_{n}$, we introduce a variable number of newlines, $n_{\text{newlines}}(l)$, sampled from a Poisson distribution $\mathcal{P}(\lambda_{\text{newlines}})$. Formally, this is expressed as:
@@ 22      Store the perturbed code as $\tilde{x}_{i}$;
25Estimate NPR: $\text{NPR}_{x}\leftarrow\text{NPR}\left(x,p_{\theta},q\right)-\frac{1}{k}\sum_{i}\text{NPR}\left(\tilde{x}_{i},p_{\theta},q\right)$;
[…]
@@ V-A Datasets
Apart from CodeSearchNet [16] described in Section III-C, we select Python code from The Stack [17] as another evaluation dataset. Similar to CodeSearchNet, The Stack provides code from a variety of open-source projects representative of real-world scenarios. We use a parsed and filtered version [42] of this dataset and also concatenate the function definitions with their corresponding comments as prompts as in [1]. For each combination of dataset and model, we sample 500 human and machine code pairs for evaluation. The maximum length of code is trimmed to 128 tokens.
@@ V-E Experimental Setup
For code generation with different models, we adopted the top-$p$ sampling strategy [36] with $p$=$0.95$ following [1]. We explored two temperature settings, $T$=$0.2$ and $T$=$1.0$, as discussed in Section III-B. The maximum length constraint for generated code was set at $128$ tokens. With respect to the perturbation-specific hyperparameters, a grid search on a held-out set from the CodeSearchNet dataset, using the SantaCoder model [3], revealed optimal values. Consequently, we set $\alpha$ and $\beta$ to $0.5$, while $\lambda_{\text{spaces}}$=3 and $\lambda_{\text{newlines}}$=2. For all experiments, we maintained a consistent configuration of generating $50$ perturbations. For the DetectGPT and DetectLLM, which involve an LLM in restoring perturbed code, we utilized the CodeT5+ (770M) model [48]. And as for the supervised baseline GPTSniffer, we trained the CodeBERT model for 5 epochs with a batch size of 16 and a learning rate of 2e-5 with AdamW optimizer. All experiments are conducted on 2 NVIDIA RTX 4090 GPUs with 24GB memory.
@@ V-F Detection Performance (RQ1)
Table IV delineates the results of various methods. According to the results, DetectCodeGPT consistently outperforms baseline methods. Compared to the strongest baseline Log Rank, our method achieves an average relative improvement of 7.6% in AUROC. In an impressive 21 of 24 combinations of dataset and model, our method provides the most accurate performance, which underscores its robustness across a variety of generative models, ranging from the 1.3 billion parameter InCoder to the 7 billion parameter CodeLlama.
We also repeated the experiments 10 times and employed a paired t-test to assess the statistical significance of the performance differences between the methods, setting the significance level at 0.05. Results show that the performance superiority of our method was statistically significant, with p-values less than 0.001. The high AUROC scores achieved across these diverse settings confirm the method’s superior capability to generalize and reliably differentiate between machine-generated and human-written code.
@@ VII Related Work
However, research on identifying machine-generated code remains relatively scarce and is purported to be more challenging than discerning machine-generated text [10]. GPTSniffer was first proposed to detect machine-generated code with supervised CodeBERT training [45]. The predominant strategies in current studies largely revolve around watermarking techniques, which embed unique markers into the code either during the training or generation [10, 58, 59]. The detection of these watermarks subsequently enables the recognition of code generated by machines. It should be noted, however, that these watermarking methods are primarily designed to address issues related to code licensing and plagiarism [60, 61]. Their reliance on modifications to the generation model renders them unsuitable for general code detection tasks.

### v5

@@ Abstract
Large language models have catalyzed an unprecedented wave in code generation. While achieving significant advances, they blur the distinctions between machine- and human-authored source code, causing integrity and authenticity issues of software artifacts. Previous methods such as DetectGPT have proven effective in discerning machine-generated texts, but they do not identify and harness the unique patterns of machine-generated code. Thus, its applicability falters when applied to code. In this paper, we carefully study the specific patterns that characterize machine- and human-authored code. Through a rigorous analysis of code attributes such as lexical diversity, conciseness, and naturalness, we expose unique patterns inherent to each source. We particularly notice that the syntactic segmentation of code is a critical factor in identifying its provenance. Based on our findings, we propose DetectCodeGPT, a novel method for detecting machine-generated code, which improves DetectGPT by capturing the distinct stylized patterns of code. Diverging from conventional techniques that depend on external LLMs for perturbations, DetectCodeGPT perturbs the code corpus by strategically inserting spaces and newlines, ensuring both efficacy and efficiency. Experiment results show that our approach significantly outperforms state-of-the-art techniques in detecting machine-generated code.  [^1] Code available at https://github.com/YerbaPage/DetectCodeGPT.
Index Terms:
machine-generated code detection, large language models, code generation, empirical analysis
@@ III-A1 Lexical Diversity
keyword def, return, else, if, for, while, $\ldots$
identifier identifier, type_identifier
literal string_content, integer, true, false, $\ldots$
operator <, >, =, ==, +, $\ldots$
syntactic symbol :, ), ], [, (, ,, ", , {, }, .
comment comment
@@ III-B Experimental Setup
We choose the state-of-the-art CodeLlama model [6] to generate code. Limited by our computational resources, we use the version with 7B parameters. As for the decoding strategies, we adopt the top-$p$ sampling method [36] with $p$=$0.95$ following [1]. The temperature $T$ is an important parameter controlling the diversity of the generated code [36]. Since current LLMs on code are usually evaluated across different decoding temperatures [5, 37, 1, 6], we generate code with $T=0.2$ and $T=1.0$ to capture the model’s behavior under different settings. The maximum length of the generated code is set to 512 tokens based on the memory constraints and the length distribution of human-written code. All experiments are conducted on 2 NVIDIA RTX 4090 GPUs with 24GB memory.
@@ III-D2 Syntax Element Distribution
The analysis of syntax element distributions, as visualized in Figure 1, reveals intriguing insights into the coding conventions and stylistic nuances between human- and machine-authored code. The “keyword”, “operator”, “syntactic symbols”, and “whitespace” proportions remain largely consistent between human and machine-authored code, across both temperatures. This consistency suggests that the foundational syntactical elements manifest similarly in both datasets. Delving deeper into the more nuanced discrepancies, a few categories emerge that underscore the differential preferences or tendencies of human and machine writers (statistical significant from Chi-square test with $p<0.01$):
@@ III-D3 Zipf’s and Heaps’ Laws
Figure 2 offers a comprehensive understanding of coding tendencies of both human and machine programmers. Starting with Zipf’s law, Figure 2(a) and 2(b) both delineate similar trends for human and machine programmers, corroborating the law’s applicability. And we can observe machine’s heightened proclivity towards tokens ranked between 10 and 100, especially at $T$=$0.2$. Turning our attention to Heaps’ law, the near-linear trends in Figure 2(c)-2(d) reaffirm the law’s validity. Also, there’s a noticeable shallowness in the slope for machine’s code at $T$=$0.2$ revealing machine’s decreased lexical diversity.
@@ III-D5 Token Likelihood and Rank
ALL -0.827 -1.658 0.831 0.319 0.811 0.492
Table VIII summarises the comparison results in terms of each token category at $T=0.2$. An intriguing finding is that whitespace tokens stand out with the highest deviation of naturalness, surpassing even the combined use of all tokens. This highlights a distinctive aspect of coding styles: machines, trained on extensive datasets, typically generate code with regular, predictable whitespace patterns. Humans, however, influenced by individual styles and practices, exhibit a wider variety in their use of whitespaces. The distinct patterns in machine-generated whitespaces, therefore, point to an inherent variation in coding style between machine and human.
@@ IV-C1 Space Insertion
Let $C$ represent the set of all possible locations to insert spaces in a code segment. We randomly select a subset $C_{s}\subseteq C$ such that $|C_{s}|=\alpha\times|C|$, where $\alpha\in[0,1]$ is a fraction representing the code locations. For each location $c\in C_{s}$, we introduce a variable number of spaces, $n_{\text{spaces}}(c)$, which is drawn from a Poisson distribution $\mathcal{P}(\lambda_{\text{spaces}})$. The Poisson distribution is chosen to simulate the randomness in human coding styles, similar to the random text infilling strategy in BART [42]. Mathematically, this can be represented as:
@@ IV-C2 Newline Insertion
We split the code into lines and obtain a set $L$ of lines. A subset $L_{n}\subseteq L$ is then chosen randomly, where $|L_{n}|=\beta\times|L|$, with $\beta\in[0,1]$ denoting the proportion of the line locations. For each line $l\in L_{n}$, we introduce a variable number of newlines, $n_{\text{newlines}}(l)$, also sampled from a Poisson distribution $\mathcal{P}(\lambda_{\text{newlines}})$:
@@ 23      Store the perturbed code as $\tilde{x}_{i}$;
26
27Estimate NPR: $\text{NPR}_{x}\leftarrow\text{NPR}\left(x,p_{\theta},q\right)-\frac{1}{k}\sum_{i}\text{NPR}\left(\tilde{x}_{i},p_{\theta},q\right)$;
28
29if $\text{NPR}_{x}>\epsilon$ then
@@ V-A Datasets
We carefully use a different split of the CodeSearchNet dataset [16] from the one used in the empirical study for evaluation. We select Python code from The Stack [17] as another evaluation dataset. Similar to CodeSearchNet, The Stack provides code from a variety of open-source projects representative of real-world scenarios. We use a parsed and filtered version [43] of this dataset and also concatenate the function definitions with their corresponding comments as prompts as in [1]. For each combination of dataset and model, we sample 500 human and machine code pairs for evaluation. The maximum length of code is trimmed to 128 tokens.
@@ V-E Experimental Setup
For code generation with different models, we adopted the top-$p$ sampling strategy [36] with $p$=$0.95$ following [1]. We explored two temperature settings, $T$=$0.2$ and $T$=$1.0$, as discussed in Section III-B. The maximum length constraint for generated code was set at $128$ tokens. With respect to the perturbation-specific hyperparameters, a grid search on a held-out set from the CodeSearchNet dataset, using the SantaCoder model [3], revealed optimal values. Consequently, we set $\alpha$ and $\beta$ to $0.5$, while $\lambda_{\text{spaces}}$=3 and $\lambda_{\text{newlines}}$=2. For all experiments, we maintained a consistent configuration of generating $50$ perturbations.
For the DetectGPT and DetectLLM, which involve an LLM in restoring perturbed code, we utilized the CodeT5+ (770M) model [49]. And as for the supervised baseline GPTSniffer, we trained the CodeBERT model for 5 epochs with a batch size of 16 and a learning rate of 2e-5 with AdamW optimizer. All experiments are conducted on 2 NVIDIA RTX 4090 GPUs with 24GB memory.
@@ V-F Detection Performance (RQ1)
Table IV delineates the results of various methods. According to the results, DetectCodeGPT consistently outperforms baseline methods. Compared to the strongest baseline Log Rank, our method achieves an average relative improvement of 7.6% in AUROC. In an impressive 20 of 24 combinations of dataset and model, our method provides the most accurate performance, which underscores its robustness across a variety of generative models, ranging from the 1.3 billion parameter InCoder to the 7 billion parameter CodeLlama.
We also repeated the experiments 10 times and employed a Wilcoxon rank sum test to assess the statistical significance of the performance differences between the methods. Results show that the performance superiority of our method was statistically significant, with p-values less than 0.001. The high AUROC scores achieved across these diverse settings confirm the method’s superior capability to generalize and reliably differentiate between machine-generated and human-written code.
@@ VII-A Machine-Generated Text Detection
VII-B Machine-Generated Code Detection
Research on identifying machine-generated code remains relatively scarce and is purported to be more challenging than discerning machine-generated text, according to the empirical study in [59]. GPTSniffer was first proposed to detect machine-generated code with supervised CodeBERT training [46]. Concurrent works [60] and [61] also explore perturbation-based methods for detecting machine-generated code, under similar framework to DetectGPT for text [9]. Our approach differs from these methods in that we perform a comprehensive empirical analysis of the differences between machine- and human-authored code. Based on the insights from the analysis, we proposed a innovative stylized perturbation strategy to achieve a more efficient and effective detection method.
Another related topic revolve around code watermarking techniques, which embed unique markers into the code either during the training or generation [10, 62, 63]. The detection of these watermarks subsequently enables the recognition of code generated by machines. It should be noted, however, that these watermarking methods are primarily designed to address issues related to code licensing and plagiarism [64, 65]. Their reliance on modifications to the generation model renders them unsuitable for general code detection tasks.
@@ VIII Conclusion
Acknowledgement
This research is supported by the National Key Research and Development Program of China (Grant No. 2023YFB4503802) and the National Natural Science Foundation of China (Grant No. 62102244). We would like to thank the anonymous reviewers for their valuable feedback and suggestions.
Appendix
Some results are omitted in the main text due to space constraints. We provide these additional results and analyses in this appendix.
I Corresponding Results for Empirical Analysis
The top tokens from machine-authored code at $T=1.0$ are shown in Table VII, corresponding to the results in Table II in the main text. Similar attention on exception handling and object-oriented programming tokens can also be observed at $T=1.0$.
Table VII: Top tokens from machine-authored code at $T=1.0$
Rank Machine-Authored Tokens
1-10 . ’ ( ) self : , = ” if
11-20 return def [ ] None in == not is import
21-30 for { raise __name__ from } - data else get
31-40 elif value path isinstance True + os ValueError key text
41-50 x type 0 1 len append replace str @ f
Figure 7 and Table VIII provide the distribution of naturalness scores and the naturalness of different categories of syntax elements, respectively. They correspond to the empirical analysis in Figure 4 and Table VIII in the main text.
We can observe from Figure 7 that the trend of naturalness scores is similar to that at $T=0.2$. although the overlap between machine- and human-authored code is greater. The naturalness of different categories of syntax elements in Table VIII shows that whitespace tokens is still the most effective feature as in $T=0.2$.
Figure 7: Distribution of naturalness scores with $T=1.0$
Table VIII: The naturalness of different categories of syntax elements with $T=1.0$
Category Log Likelihood Log Rank
Machine Human $\Delta$ Machine Human $\Delta$
keyword -2.025 -2.128 0.103 0.968 1.053 0.085
identifier -0.787 -0.874 0.087 0.328 0.378 0.050
literal -1.059 -1.364 0.305 0.454 0.630 0.176
operator -1.614 -1.835 0.221 0.733 0.872 0.139
symbol -0.968 -1.639 0.671 0.321 0.781 0.460
comment -2.395 -3.028 0.633 1.180 1.610 0.430
whitespace -2.058 -2.740 0.682 0.946 1.441 0.495
ALL -1.367 -1.658 0.291 0.618 0.811 0.193
II Cross-model Detection Performance at $T=1.0$
Figure 8 illustrates the cross-model detection performance at $T=1.0$, corresponding to the results in Figure 6 in the main text. The results show that the cross-model detection performance of DetectCodeGPT has a similar trend across different temperatures.
Figure 8: Cross-model detection performance at $T=1.0$
III Impact of Varying Length for Detection
The maximum length of code is trimmed to 128 tokens in our experiments. Figure 9 illustrates the trend in detection performance as the length of code snippets varies across different temperatures, compared against two of the most competitive baselines. This experiment uses the Stack dataset with CodeLlama. The results show that trimming the first 128 tokens of the snippets is sufficient to maintain high detection performance.
(a) AUROC at $T=0.2$
(b) AUROC at $T=1.0$
Figure 9: AUROC with different code trimming length